# LLMs as Hackers: Autonomous Linux Privilege Escalation Attacks

### v4

@@
Abstract.
Penetration testing, an essential component of software security testing, allows organizations to identify and remediate vulnerabilities in their systems, thus bolstering their defense mechanisms against cyberattacks. One recent advancement in the realm of penetration testing is the utilization of Language Models (LLMs). We explore the intersection of LLMs and penetration testing to gain insight into their capabilities and challenges in the context of privilege escalation. We introduce a fully automated privilege-escalation tool designed for evaluating the efficacy of LLMs for (ethical) hacking, executing benchmarks using multiple LLMs, and investigating their respective results.
Our results show that GPT-4-turbo is well suited to exploit vulnerabilities (33–83% of vulnerabilities). GPT-3.5-turbo can abuse 16–50% of vulnerabilities, while local models, such as Llama3, can only exploit between 0 and 33% of the vulnerabilities.
We analyze the impact of different context sizes, in-context learning, optional high-level guidance mechanisms, and memory management techniques. We discuss challenging areas for LLMs, including maintaining focus during testing, coping with errors, and finally comparing LLMs with human hackers.
The current version of the LLM-guided privilege-escalation prototype can be found at https://github.com/ipa-labs/hackingBuddyGPT.
1. Introduction
In the rapidly evolving field of cybersecurity, penetration testing (“pen-testing” or “hacking”) plays a pivotal role in identifying and mitigating potential vulnerabilities. A crucial subtask of pen-testing is privilege escalation, which involves exploiting a bug, design flaw, or configuration oversight in an operating system or software application to gain elevated access to resources that are normally protected from an application or user (Wikipedia, [n. d.]). The ability to escalate privileges provides a malicious actor with increased access, potentially leading to more significant breaches or system damage. Therefore, understanding and improving the performance of tools used for this task is highly relevant.
In this paper, we focus on investigating the performance of Large Language Models (LLMs) in the context of penetration testing, specifically for Linux privilege escalation. LLMs have shown remarkable abilities to emulate human behavior that can be used to automate and improve various tasks in pen-testing (Deng et al., 2023; Happe and Jürgen, 2023). However, there is currently no understanding on how these models perform in common privilege escalation scenarios.
@@ 1. Introduction
  • an fully-automated LLM-driven Linux privilege escalation prototype, wintermute (Section 3 Methodology)
@@
2. Background and Related Work
The background section focuses on the two distinct areas that this work integrates: LLMs and privilege escalation.
@@ 2. Background and Related Work
Training an LLM incurs high costs. Recently, alternative approaches have tried to achieve high performance while avoiding expensive training. In-Context Learning (Dong et al., 2022; Dai et al., 2023) includes background information within the prompt and thus exchanges knowledge inherently stored within the model with external knowledge. Similarly, Chain-of-Thought prompting includes step-by-step answer examples within the context (Kojima et al., 2022). Both approaches make the context a very limited resource.
Real-world tasks often must be split into smaller subtasks or steps. Multiple approaches try to emulate this through LLMs, ranging from minimal approaches such as BabyAGI (Nakajima, [n. d.]) to Tree-of-Thoughts (Yao et al., 2023) or Task-Lists (Deng et al., 2023). Our prototype utilizes an approach similar to BabyAGI’s minimal approach.
A combination of the mentioned topics, i.e., small viable model sizes, using context for adding information while having enough context to describe the task at hand and having task/state management for keeping track of sophisticated work, would make LLMs viable for local usage or for usage with private/sensitive data.
@@
LLM benchmarks are typically based on common sense reasoning tasks. This is sensible, as common-sense reasoning is a transferable skill well suited to many tasks, including penetration-testing. However, a recent survey by Davis (Davis, 2023) shows that many existing common sense reasoning benchmarks have quality issues within their tasks. Another open question is if high scores in synthetic common-sense benchmarks translate into high scores in real-world domain-specific scenarios, which are typically not tested by LLM makers.
2.1. LLM usage by Black-/White-Hat Hackers
The potential of LLMs is seen by ethical hackers and by not-so-legal ones (Black-Hats). Gupta et al. identify multiple areas of interest for using LLMs (Gupta et al., 2023) including phishing/social engineering, pen-testing and the generation of malicious code/binaries, be it payloads, ransomware, malware, etc.
Recent darknet monitoring (Gatlan, 2023) indicates that Black-Hats are already offering paid-for LLMs: One suspected threat actor is offering WormGPT (Mascellino, 2023) and FraudGPT: while the former focuses upon social engineering, the latter aids writing malicious code, malware, payloads. The same threat actor is currently preparing DarkBert (Montalbano, 2023)l which is supposedly based on the identically named DarkBERT (Jin et al., 2023), a LLM that was designed to combat cybercrime. Other darknet vendors also offer similar products: XXXGPT is advertised for malicious code creation, WolfGPT is advertised for social engineering (Dutta, 2023). Please note that all those products are offered within the darknet behind paywalls, so their claims cannot be independently verified.
2.1.1. Hacking with LLMs
To the best of our knowledge, there is currently no darknet-offered LLM-aided penetration testing tool. But, as the other areas have shown, this is just a matter of time.
pentestGPT utilizes LLMs for solving CTF-style challenges (Deng et al., 2023). It is an interactive tool that guides pen-testers both on a high-level (pen-testing approach) and on a low level (tool selection and execution). It employs a hierarchical state model to keep track of the current penetration testing progress. Their github repository explicitly recommends using GPT-4 over GPT-3.5 as the latter “leads to failed tests in simple tasks”. Compared to pentestGPT, our prototype focuses upon fully automated penetration-testing without interactive user feedback to allow for automated benchmarks. In addition, we tested local LLMs for their feasibility for pen-testing.
pentestGPT uses HackTheBox cloud-based challenges for their evaluation. To allow for greater control, our chosen benchmark is based upon locally generated and operated virtual machines. By narrowing the scope to Linux privilege-escalation vulnerabilities, we are able to more deeply analyze the differences between the different LLMs hoping that future research can base their model selection upon firmer foundations.
2.2. Linux Privilege-Escalation Vulnerabilities
Table 1. Benchmark Test-Scenarios (Happe and Cito, 2024)
Vulnerability-Class Name Description
SUID/sudo files suid-gtfo exploiting suid binaries
SUID/sudo files sudo-all sudoers allows execution of any command
SUID/sudo files sudo-gtfo GTFO-bin in sudoers file
priv. groups/docker docker user is in docker group, docker is running in privileged mode
password hygiene password reuse root uses the same password as lowpriv
password hygiene weak password root is using the password “root”
password hygiene password in user text file vacation.txt in the user’s home directory with the root password
information disclosure password in user config file reused password in the local database configuration
information disclosure bash_history root password is in .bash_history
information disclosure SSH key lowpriv can use key-bases SSH without password to become root
cron-based cron file with write access is called through cron as root
cron-based cron-wildcard cron backups the backup directory using wildcards
@@ 2.2. Linux Privilege-Escalation Vulnerabilities
Privilege-Escalation (short priv-esc) is the art of making a system perform operations that the current user should not be allowed to. We focus upon a subsection of priv-esc, namely local Linux low-privilege users trying to become root (uid 0, the all powerful system administrator). This is a common task that occurs after an initial system breach.
2.2.1. Benchmarks
Due to the sensitive, unpredictable, and potentially destructive nature of our experiments, the security and safety of our test environment is of high importance, i.e., the commands given by the LLM to be executed within our test environment must not interact with any non-test system nor network. To achieve this, the test scenarios should be hosted within virtual machines upon a virtual network that is not publicly reachable.
There is a lack of offline Linux Priv-Esc benchmarks. We assume that one of the reasons is the competitive nature of security testing: as soon as a benchmark is established, tools can optimize for their test-cases, and thus invalidate the benchmark. In contrast, ample unstructured information about Linux privilege escalation techniques can be found in public online wikis such as hacktricks [^1] https://book.hacktricks.xyz/ or GTFObins [^2] https://gtfobins.github.io/, a collection of privilege escalation techniques.
2.2.2. Automated “Hacking”
@@
Kowira et al. give an overview of existing Linux enumeration scripts and state the lack of automated Linux privilege escalation scripts (Kowira et al., 2024). Penetration testers have to parse the various enumeration scripts’ outputs and match the provided information with potential exploitation attacks. In contrast, we instruct LLMs to autonomously enumerate and execute privilege-escalation attacks.

### v5

@@
Abstract
Penetration testing allows organizations to identify and remediate vulnerabilities in their systems. It is an essential part of active software security testing as it bolsters defense mechanisms against cyberattacks. One recent advancement in the realm of penetration testing is the utilization of Language Models (LLMs). We explore the intersection of LLMs and penetration testing to gain insight into their capabilities and challenges, esp. in the context of privilege escalation.
We introduce a fully automated privilege-escalation tool designed for evaluating the efficacy of LLMs for ethical hacking, executing benchmarks using multiple LLMs, and investigating their respective results.
A localized benchmark allows us to evaluate the effectiveness of privilege-escalation techniques performed by both humans and automated tooling. To address this gap, we developed a comprehensive benchmark for Linux privilege escalation. It provides a standardized platform to evaluate and compare the performance of human and synthetic actors, e.g., hacking scripts or automated tooling in the form of an agentic LLM system.
We analyze the impact of different context sizes, in-context learning, high-level guidance mechanisms, and memory management techniques. We discuss challenging areas for LLMs, including maintaining focus during testing, coping with errors, and finally comparing LLMs with human hackers. Our results show that GPT-4-turbo is well suited to exploit vulnerabilities (33–83% of vulnerabilities). GPT-3.5-turbo can abuse 16–50% of vulnerabilities, while local models, such as Llama3, can only exploit between 0 and 33% of the vulnerabilities.
The current version of the LLM-guided privilege-escalation prototype can be found at https://github.com/ipa-labs/hackingBuddyGPT. Our Linux Privilege Escalation benchmark can be retrieved from https://github.com/ipa-labs/benchmark-privesc-linux. Both are already in use by third parties.
1 Introduction
In the rapidly evolving field of cybersecurity, penetration testing (“pen-testing” or “hacking”) plays a pivotal role in identifying and mitigating potential vulnerabilities. A crucial subtask of pen-testing is privilege escalation, which involves exploiting a bug, design flaw, or configuration oversight in an operating system or software application to gain elevated access to resources that are normally protected from an application or user [^1] https://en.wikipedia.org/wiki/Privilege_escalation. The ability to escalate privileges provides a malicious actor with increased access, potentially leading to more significant breaches or system damage. Therefore, understanding and improving the performance of tools used for this task is highly relevant and impacts real-life security.
In this paper, we focus on investigating the performance of Large Language Models (LLMs) in the context of penetration testing, specifically for Linux privilege escalation. LLMs have shown remarkable abilities to emulate human behavior that can be used to automate and improve various tasks in pen-testing (Deng et al., 2023; Happe and Cito, 2023a). However, there is currently no understanding on how these models perform in common privilege escalation scenarios.
@@ 1 Introduction
  • a publicly available Linux Privilege Escalation benchmark set that can be run on local premises due to the safety- and security-critical nature of this benchmark (Section 4).
  • an fully-automated LLM-driven Linux privilege escalation prototype (Section 3 Methodology). We release the prototype including all prompts on github as open-source.
@@
While these detail our high-level goals, some of the questions we’ve answered within this work are:
  • We show that autonomous penetration-testing using a single LLM-callout control loop is a resource-efficient alternative to commonly used multi-LLM agent based systems. Compared to state-of-the-art autonomous multi-LLM solutions which achieve 7–20% hacking success rates, we achieve up to 66% in unguided scenarios. This is comparable to a human baseline of 75%.
  • When adding high-level guidance, similar to the guidance provided by human penetration testers or the Planner module of pentestGPT-inspired architectures, the success rate rises to 83%. This is comparable to the success rate of our human baseline of 91%.
  • The state of the art offers conflicting guidance about the efficacy of adding background hacking information as well as on the impact of using larger execution histories. This work investigates these questions empirically, indicating that background hacking information has a smaller impact when using state-of-the-art models such as GPT-4 while increasing the stored history increased our success rates significantly.
2 Background and Related Work
The background section focuses on the two distinct areas that this work integrates: LLMs and penetration testing, also known as “hacking”.
2.1 Large-Language Models
@@
An alternative to using small language models is quantizing models with larger parameter counts. In this approach, parts of the model weights are quantized from 32bit floating points into data types of lower precision, e.g., 4 bit integers. This reduces the model’s memory requirements, and thus makes local model usage computationally feasible. There is an ongoing discussion on the relative trade-off between using smaller models of full precision and using larger quantized models (Huang et al., 2024).
Training an LLM incurs high costs. Recently, alternative approaches have tried to achieve high performance while avoiding expensive training. In-Context Learning (Dong et al., 2022; Dai et al., 2023) includes background information within the prompt and thus exchanges knowledge inherently stored within the model with external knowledge. Similarly, Chain-of-Thought prompting includes step-by-step answer examples within the context (Kojima et al., 2022). Both approaches make the context a very limited resource, even with ever increasing context-sizes of newer model generations.
Real-world tasks often must be split into smaller subtasks or steps. Multiple approaches try to emulate this through LLMs, ranging from minimal approaches such as BabyAGI (Nakajima, 2023) to Tree-of-Thoughts (Yao et al., 2023) or Task-Lists (Deng et al., 2023). Wang et al. (2023) name this approach in a contemporary paper plan-and-solve.
A combination of the mentioned topics, i.e., small viable model sizes, using context for adding information while having enough context to describe the task at hand and having task/state management for keeping track of sophisticated work, would make LLMs viable for local usage, esp. with sensitive locally-stored data.
@@
LLM Benchmarks. LLM benchmarks are typically based on common sense reasoning tasks. This is sensible, as common-sense reasoning is a transferable skill well suited to many tasks, including penetration-testing. However, a recent survey by Davis (2023) shows that many existing common sense reasoning benchmarks have quality issues within their tasks. Another open question is if high scores in synthetic common-sense benchmarks translate into high scores in real-world domain-specific scenarios, which are typically not tested by LLM makers.
2.2 Penetration Testing
Penetration Testing, short pen-testing, is described by Geer and Harthorne (2002) as “the art of finding an open door”. Its goal is to find a vulnerability within the subject-under-test to falsify the hypothesis that the subject is secure. The outcome of a penetration test allows defenders to fortify their systems so that other, potentially malicious, attackers cannot abuse similar vulnerabilities (Bishop, 2007). Professionals performing those tests are typically called penetration-testers, pen-tester, or simply hackers. An additional differentiation is often performed upon the intend of the pen-tester: white hats perform ethical research to improve the field of software security while black-hats are malicious and work for monetary or political gain.
Shah and Mehtre (2015) further elaborate on the nature of penetration testing and differentiate between Vulnerability Assessments and Penetration Testing. The goal of the former is to identify as many possible vulnerabilities within the subject-under-test as possible, while the latter emulates an attacker that tries to actively exploit a found vulnerability. As Penetration Testing can lead to system instabilities and data loss, automated tooling often focus upon well-known Vulnerability Assessment, not exploitation (Shebli and Beheshti, 2018). Tooling such as nmap [^2] https://nmap.org/, OpenVAS [^3] https://www.openvas.org/, PortSwigger BURP [^4] https://portswigger.net/burp or ZAP [^5] https://www.zaproxy.org/ often utilize rule-based detection systems as well as databases of known vulnerable software versions and use more-aggressive techniques such as fuzzing only upon explicit user interaction and as measures of last resort.
Only little empirical research into how penetration testers perform their work, and the potential problems therein, has been performed. Happe and Cito (2023b) performed an interview study with professional penetration testers. One of their key findings was that security researchers and security practitioners (penetration testers) differ in their methodologies and tooling. While security researchers focus upon finding new and novel vulnerabilities, i.e., finding 0-days, security practitioners spend the majority of their time using known vulnerabilities and abusing security misconfigurations. These are often emulated through Capture-the-Flag (CTF) challenges which enable transfer learning. When attacking enterprise networks or performing privilege-escalation attacks, interviewees mentioned that they would never search for novel 0-day vulnerabilities due to their limited amount of time, but depend upon their knowledge of existing vulnerabilities as detailed by large online knowledge bases such as hacktricks [^6] https://book.hacktricks.xyz/.
@@ 2.2 Penetration Testing
Linux Privilege-Escalation Vulnerabilities. Privilege-Escalation (short priv-esc) is the art of making a system perform operations that the current user should not be allowed to. We focus upon a subsection of priv-esc, namely local Linux low-privilege users trying to become the all powerful root system administrator indicated by an user id of 0. This is a common task that occurs after an initial system breach.
Privilege-Escalation attacks are typically performed manually by searching for exploitable configurations or vulnerable tools. The initial act of system reconnaissance, often named enumeration, is often automated through usage of tools such as linpeas.sh [^7] https://github.com/peass-ng/PEASS-ng/tree/master/linPEAS. Exploitation itself is typically done manually through the, hopefully ethical, hacker.
In many security areas, established standards and methodologies guide novice practitioners, e.g., in the web application area the non-profit organisation OWASP provides both the de-facto standard list of commonly used web vulnerabilities [^8] https://owasp.org/Top10/ as well as detailed testing guides [^9] https://owasp.org/www-project-web-security-testing-guide/stable/. In contrast, there is no such coverage for Linux Privilege Escalation Attacks. Partially fitting is the MITRE ATT&CK framework [^10] https://attack.mitre.org/resources/faq/#other-models-faq that “is a knowledge base of cyber adversary behavior and taxonomy for adversarial actions across their lifecycle”. Originally focusing upon Microsoft Windows Enterprise networks, subsequent iterations also include Linux attack vectors. MITRE ATT&CK does not offer a methodology, i.e., it does not describe attacks paths, but is an unordered taxonomy of potential attack vectors, thus does not provide high-level guidance to security practitioners, nor can be used as a high-level structure for benchmarks.
Instead of established standards, aspiring penetration testers typically consume living online information sources. Ample unstructured information about Linux privilege escalation techniques can be found in public online wikis such as hacktricks [^11] https://book.hacktricks.xyz/ or GTFObins [^12] https://gtfobins.github.io/, a collection of privilege escalation techniques. In addition, Capture-the-Flag (CTF) style exercises allow penetration testers to hone their skills. Sites such as TryHackMe [^13] https://tryhackme.com/ or HackTheBox [^14] https://www.hackthebox.com/ allow online access to an ever-changing set of vulnerable virtual machines.
@@
Automated Linux Priv-Esc Tools. Kowira et al. (2024) give an overview of existing Linux enumeration scripts and state the lack of automated Linux privilege escalation. Penetration testers have to parse the various enumeration scripts’ outputs and match the provided information with potential exploitation attacks. In contrast, we investigate the usage of LLMs to autonomously enumerate and execute privilege-escalation attacks.